Help step by step please1. For each of the following boxes, write the corresponding random variable and its distribution. The random variable, call i
Answer #1

As corresponding probability for X are not given hence we will use equal probabilities for each cases.

a) There are 4 values hence for each the probability is 1/4

x 0 1 1 6
p 1/4 1/4 1/4 1/4

\small E(X)=\sum xp=0*\frac{1}{4}+1*\frac{1}{4}+1*\frac{1}{4}+6*\frac{1}{4}=\frac{8}{4}=2

Hence E(x) = 2 for part a)

And expected values of 100 draws = 100*E(x) = 100*2 = 200

b)

x -2 -1 0 2
p 1/4 1/4 1/4 1/4

\small E(X)=\sum xp=-2*\frac{1}{4}-1*\frac{1}{4}+0*\frac{1}{4}+2*\frac{1}{4}=\frac{-1}{4}=-0.25

Hence E(x) = -0.25 for part b)

And expected value of 100 draws = 100*E(x) = 100* (-0.25) = -25

Please comment if any doubt. thank you.